OpenStudy (anonymous):

What is the value of f[g(3)] for the functions f(x) = x + 2 and g(x) = 5x - 1? (Type only the value of your final answer, no variables.) Numerical Answers Expected!

OpenStudy (kainui):

Too bad!

OpenStudy (ghazi):

first of all find f(g(x)) put x=5x-1 in the function F(x) to find f(g(x))

OpenStudy (ghazi):

i.e f(g(x))=(5x-1)+2

OpenStudy (anonymous):

\[g(3)=5\times 3-1=15-1=14\]now find \[f(14)\]

OpenStudy (anonymous):

\[f(14)=14+2=16\]